# Identifying Factual Inconsistencies in Summaries: Grounding Model Inference via Task Taxonomy

### v2

@@ Abstract
Factual inconsistencies pose a significant hurdle for the faithful summarization by generative models. While a major direction to enhance inconsistency detection is to derive stronger Natural Language Inference (NLI) models, we propose an orthogonal aspect that underscores the importance of incorporating task-specific taxonomy into the inference. To this end, we consolidate key error types of inconsistent facts in summaries, and incorporate them to facilitate both the zero-shot and supervised paradigms of LLMs. Extensive experiments on ten datasets of five distinct domains suggest that, zero-shot LLM inference could benefit from the explicit solution space depicted by the error type taxonomy, and achieves state-of-the-art performance overall, surpassing specialized non-LLM baselines, as well as recent LLM baselines. We further distill models that fuse the taxonomy into parameters through our designed prompt completions and supervised training strategies, efficiently substituting state-of-the-art zero-shot inference with much larger LLMs.
@@
Grounding Model Inference via Task Taxonomy
Liyan Xu1thanks: Equal contributions.  Zhenlin Su211footnotemark: 1  Mo Yu1  Jin Xu2,3  Jinho D. Choi4  Jie Zhou1  Fei Liu4 1Pattern Recognition Center, WeChat AI 2South China University of Technology 3Pazhou Lab, Guangzhou  4Emory University liyanlxu@tencent.comjinxu@scut.edu.cnfei.liu@emory.edu
@@ 1 Introduction
While it is a promising direction to keep enhancing factual inconsistency recognition by deriving stronger NLI models, such as FactCC Kryscinski et al. (2020), DocNLI Yin et al. (2021), FalseSum Utama et al. (2022), AMRFact Qiu et al. (2024), in this work, we propose approaches from an orthogonal aspect, which examines the incorporation of explicit solution space into the inference, such that either zero-shot prompting or a trained model reaches decisions according to explicit task-specific cues, i.e. an explicit error type taxonomy.
@@
Our motivation stems from the distinct nature between summary inconsistencies and NLI: summaries are grounded by the original document, thus leaning towards reiteration, whereas NLI tackles a broader problem that involves extrapolation. Since we roughly deem the scope of summary inconsistency detection as smaller, one can consolidate and leverage its task-specific taxonomy to rationalize a more effective and explainable inference.
@@
As there exist numerous annotation schemas adopted by previously introduced datasets, factual errors are firstly unified into a fine-grained taxonomy by AggreFact Tang et al. (2023), which we consolidate upon and identify five common error types that are salient for recognizing summary inconsistencies, including Predicate Error, Entity Error, Circumstantial Error, Coreference Error and Addition Error (Section 3), covering a wide variety of datasets (Table 1). The identified error types are then utilized to anchor the inference of factual inconsistencies. Specifically, we examine their efficacy with LLMs in both zero-shot and supervised paradigms, and demonstrate the utility of task-specific taxonomy in complementary to the sole NLI-style classification.
@@
For the zero-shot setting (Section 4), we craft the instruction tailored for each error type in the prompt, directing LLMs to reason specific error types according to the given guidance. To handle long summaries, we additionally propose a window-based prompting scheme, as an effective alternative to the vanilla prompting. For a comprehensive evaluation, our experiments are conducted on 10 datasets across five domains, including summarization on different news sources, daily or professional dialogues, official reports and narrative stories. Moreover, we employ models from OpenAI (ChatGPT, GPT-4o) along with strong open-source LLMs including Llama-3 Touvron et al. (2023) and Mistral Jiang et al. (2023) to ensure a robust conclusion. Empirical results suggest that our proposed methods surpass all baselines, including 7 non-LLM baselines and 4 LLM baselines, showing that zero-shot LLM inference could benefit from a grounded solution space by depicting the task taxonomy in the instruction. Our proposed methods, termed Factuality with Taxonomy (FacTax), achieve the best overall performance across five domains; especially, FacTax with ChatGPT outperforms previous state-of-the-art zero-shot baseline G-Eval with GPT-4 Liu et al. (2023a); Qiu et al. (2024) on the AggreFact-FtSota benchmark Tang et al. (2023).
@@
We then further seek to distill a model that fuses the task taxonomy into model parameters through supervised training. By unifying the error types of previous independently introduced datasets, we regard them jointly as training resources. Llama3-8B models are trained to learn binary decisions as well as to recognize specific error types on summaries, through our designed completions and training strategies. The resulting trained model outperforms previous supervised baselines, and is able to match the best zero-shot inference performance, effectively acting as an efficient alternative to zero-shot reasoning with much larger LLMs.
@@
  • We pinpoint key error types and incorporate them into our designed zero-shot prompting schemes, anchoring LLM reasoning within an explicit solution space. Experiments on diverse datasets and LLMs demonstrate its efficacy over baselines.
  • We further distill a model that rationalizes the task taxonomy into parameters through our supervised training strategies, practically substituting zero-shot reasoning with SOTA performance.
@@ 2 Related Work
Factual Inconsistency Evaluation Datasets
Non-LLM Approaches
State-of-the-art models prior to LLMs mainly focus around two directions. The first is to effectively leverage NLI models to assess the entailment between the document-summary pair, such as Falke et al. (2019) and SummaC Laban et al. (2022). Within this direction, several works focus on improving the NLI model itself through methods such as synthetic data construction Kryscinski et al. (2020); Yin et al. (2021); Utama et al. (2022); Qiu et al. (2024) or multitask learning Zha et al. (2023a, b). The second direction employs QA-based models, such as QuestEval Scialom et al. (2021) and QAFactEval Fabbri et al. (2022), where they generate questions regarding explicit entities in the summary, then verify upon the source document. Besides the two main directions, other works have also explored to recognize factual errors through methods such as syntactic dependencies Goyal and Durrett (2020) or information extraction Nan et al. (2021).
LLM Approaches
The capability of LLMs on detecting inconsistencies have been studied by several recent works. Most of them resolve this task in the zero-shot or few-shot prompting manner Shen et al. (2023); Luo et al. (2023); Wang et al. (2023); Liu et al. (2023a). Other utilization of LLMs have also been proposed, such as synthetic data generation with LLMs Gekhman et al. (2023).
3 Task Taxonomy
@@
Table 1 shows the resulting conversion as well as more statistical details. We do notice that neither our adopted taxonomy nor the fine-grained one in AggreFact is completely free from noises, due to different annotation standards across datasets.
@@ 4 Approach: Zero-Shot Paradigm
To incorporate the error type taxonomy, we first propose zero-shot prompting methods that leverage the general comprehension capability of LLMs, aiming to depict the explicit solution space to facilitate the zero-shot inference.
@@ 4.1 FacTax
Our first designed prompting scheme, dubbed Factuality with Taxonomy (FacTax), follows the standard zero-shot procedure: for a document-summary pair, we instruct a LLM to determine whether the summary is factually correct, as in previous works utilizing LLMs Luo et al. (2023); Wang et al. (2023); Liu et al. (2023a). For each error type, we handcraft its explanation along with an optional example, and we ask the LLM to reason in a Chain-of-Thought (CoT) style Wei et al. (2022): whether there are any specific error types present in the summary, instead of generating a binary decision directly. A summary is thus recognized as factually correct through a rationalization stage, when no specified error types are present.
The resulting zero-shot inference to this end, is regularized by the underlying task taxonomy, so to achieve a comprehensive task reasoning. We provide our full prompt in Appx. B.
@@ 4.2 FacTax-WD
After our preliminary experiments, we manually set the window size as roughly 30 words to balance between the performance and efficiency. It is worth noting that smaller window size (i.e. one sentence per window) does not necessarily lead to higher performance, as we observe that pronouns in the summary could often introduce false negatives when inferred without its broader context.
@@
Table 2: Evaluation results for the zero-shot paradigm (Section 4.3). Five domains (10 datasets in total) are evaluated, where the setting for CNN/DM and XSum is kept consistent and comparable with AggreFact-FtSota Tang et al. (2023) using thresholds per dataset. MACRO is the final evaluation metric that computes the macro-average score across each domain. FacTax methods are our proposed approaches that ground the zero-shot inference by the task taxonomy. All LLM-based methods are shown the averaged scores of three repeated runs for robust evaluation, and they are directly comparable due to adopting the same underlying model (gpt-3.5-turbo-0125).
@@ 4.3 Zero-Shot Experiments
For GovReport and SQuALITY, documents are long articles that exceed certain models’ length limit. We follow Wu et al. (2023) that for each document, top sentences that maximize ROUGE scores towards the summary are retrieved as a condensed context (details in Appx. C).
@@
We adopt strong non-LLM models as baselines, including QuestEval Scialom et al. (2021), QAFactEval Fabbri et al. (2022), SummaC Laban et al. (2022), AlignScore Zha et al. (2023a) and Align Zha et al. (2023b). For each, we either take the evaluation scores from prior works, or run the code released by the authors on new datasets not evaluated previously. Additionally, we also include scores of Falsesum and AMRFact from their original papers. More details on non-LLM baselines are provided in Appx. C.
@@
We use ChatGPT-ZS, ChatGPT-CoT Luo et al. (2023), ChatGPT-Star Wang et al. (2023), and G-Eval Liu et al. (2023a) as the LLM baselines, which have achieved strong performance in AggreFact. For direct comparison, we use the same ChatGPT (gpt-3.5-turbo-0125) for all prompts. In later analysis (Sec. 4.4), we vary LLMs for more insights on model comparison.
For FacTax, we adapt the prompts to additionally yield a score on GovReport and SQuALITY summaries to enable evaluation with their score-based labels.
@@ Results
Due to the variation from sampled generation of LLMs, we run all LLM-based methods three times for robust conclusions, and show the averaged scores of each dataset in Table 2, along with evaluation results of non-LLM baselines. Several observations from Table 2 can be made as follows.
@@ 4.4 Zero-Shot Analysis
We focus on three common domains with shorter summaries: CNN/DM, XSum, Dialogues, and perform further analysis for more regarding insights.
@@
Table 3: Evaluation results using FacTax on three domains by varying LLMs. FacTax with ChatGPT achieves SOTA performance on AggreFact-FtSota benchmark, surpassing previous zero-shot SOTA G-Eval with GPT-4 (from Qiu et al. (2024)).
@@
Apart from ChatGPT, we also employ GPT-4o from OpenAI, as well as strong open-source LLMs including Llama3-8B/70B and Mistral 8B. Table 3 provides the model comparison by adopting the same FacTax prompts. Notably, FacTax with ChatGPT achieves SOTA performance on AggreFact-FtSota benchmark Tang et al. (2023), outperforming G-Eval with GPT-4 Qiu et al. (2024).
Just by switching to GPT-4o, there comes a direct boost upon ChatGPT by 3 MACRO score overall. There is still quite a gap of almost 10% between the smaller 7B/8B models and the larger OpenAI models. By Table 3, it is evident that increasing the model size significantly improves the task reasoning, as switching Llama3-8B to 70B obtains performance gain by an impressive 13.6%.
@@
Figure 2 plots the performance curve on different lengths of summaries using ChatGPT. FacTax-WD is shown more robust against long summaries, due to its length-agnostic scoring mechanism. For long summaries, false positives become more often for FacTax, which is alleviated by the window-based inference of FacTax-WD.
@@ 5 Approach: Supervised Paradigm
As Section 4 has demonstrated the strengths of grounding zero-shot LLM inference by its task taxonomy, we further seek to distill a model that absorbs the taxonomy into LLM parameters through supervised training. Two advantages could come with such distillation. First, by learning the taxonomy from examples, the model gains real-world distribution of each error type, rather than relying on shallow comprehension through prompt instructions. Second, it is more efficient and practical compared to zero-shot methods, avoiding the need for large model sizes and lengthy generation due to zero-shot CoT reasoning.
@@ Results
Table 4 shows the evaluation results of our supervised paradigm, along with comparison by various zero-shot results. Unsurprisingly, trained Llama3-8B models of any settings outperform its zero-shot inference by large margins, up to 13 MACRO score. More importantly, I&II-Taxonomy + INF-Taxonomy achieves the best performance, matching the best existing approaches by using FacTax with GPT-4o and Llama3-70B. Our trained model can effectively serve as an efficient alternative to zero-shot inference by much larger LLMs, which we will publicly release for the research community.

### v3

# Identifying Factual Inconsistencies in Summaries: Grounding LLM Inference via Task Taxonomy

@@ Abstract
Factual inconsistencies pose a significant hurdle for the faithful summarization by generative models. While a major direction to enhance inconsistency detection is to derive stronger Natural Language Inference (NLI) models, we propose an orthogonal aspect that underscores the importance of incorporating task-specific taxonomy into the inference. To this end, we consolidate key error types of inconsistent facts in summaries, and incorporate them to facilitate both the zero-shot and supervised paradigms of LLMs. Extensive experiments on ten datasets of five distinct domains suggest that, zero-shot LLM inference could benefit from the explicit solution space depicted by the error type taxonomy, and achieves state-of-the-art performance overall, surpassing specialized non-LLM baselines, as well as recent LLM baselines. We further distill models that fuse the taxonomy into parameters through our designed prompt completions and supervised training strategies, efficiently substituting state-of-the-art zero-shot inference with much larger LLMs. Our data and code are publicly released at https://github.com/lxucs/factax.
@@
Grounding LLM Inference via Task Taxonomy
Liyan Xu1thanks: Equal contributions.  Zhenlin Su211footnotemark: 1  Mo Yu1  Jin Xu2,3thanks: Co-corresponding authors.  Jinho D. Choi4  Jie Zhou1  Fei Liu422footnotemark: 2 1Pattern Recognition Center, WeChat AI 2South China University of Technology 3Pazhou Lab, Guangzhou  4Emory University liyanlxu@tencent.comzhenlinsu75@gmail.comjinxu@scut.edu.cnfei.liu@emory.edu
@@ 1 Introduction
While it is a promising direction to keep enhancing factual inconsistency recognition by deriving stronger NLI models, such as FactCC Kryscinski et al. (2020), DocNLI Yin et al. (2021), FalseSum Utama et al. (2022), AMRFact Qiu et al. (2024), in this work, we propose approaches from an orthogonal aspect, which examines the incorporation of explicit solution space into the inference, such that either zero-shot LLM prompting or trained models are grounded by explicit task-specific cues, i.e. an explicit error type taxonomy.
@@
Our motivation stems from the distinct nature between summary inconsistencies and NLI: summaries are grounded by the original document, thus leaning towards reiteration, whereas NLI tackles a broader problem that involves extrapolation. Since the scope of summary inconsistency detection is roughly smaller than NLI, one can consolidate and leverage its task-specific taxonomy to rationalize a more effective inference with interpretability.
@@
As there exist numerous annotation schemas adopted by previously introduced datasets, factual errors have been unified into a fine-grained taxonomy by AggreFact Tang et al. (2023), which we consolidate upon and identify five common error types that are salient for recognizing summary inconsistencies, including Predicate Error, Entity Error, Circumstantial Error, Coreference Error and Addition Error (Section 3), covering a wide variety of datasets (Table 1). The identified error types are then utilized to anchor the inference of factual inconsistencies. Specifically, we examine their efficacy with LLMs in both zero-shot and supervised paradigms, and demonstrate the utility of task-specific taxonomy in complementary to the sole NLI-style classification.
@@
For the zero-shot setting (Section 4), we craft the instruction tailored for each error type in the prompt, directing LLMs to reason specific error types according to the given guidance. To handle long summaries, we additionally propose a window-based prompting scheme, as an effective alternative to the vanilla prompting. For a comprehensive evaluation, our experiments are conducted on 10 datasets across five domains, including summarization on different news sources, daily or professional dialogues, official reports and narrative stories. Moreover, we employ models from OpenAI (ChatGPT, GPT-4o) along with strong open-source LLMs including Llama-3 Touvron et al. (2023) and Mistral Jiang et al. (2023) towards a robust conclusion.
Empirical results suggest that our proposed methods surpass all baselines, including 7 non-LLM baselines and 4 LLM baselines, showing that zero-shot LLM inference could benefit from a grounded solution space by depicting the task taxonomy in the instruction. Our proposed methods, termed Factuality with Taxonomy (FacTax), achieve the best overall performance across five domains; especially, FacTax with ChatGPT outperforms previous state-of-the-art zero-shot baseline G-Eval with GPT-4 Liu et al. (2023a); Qiu et al. (2024) on the AggreFact-FtSota benchmark Tang et al. (2023). Our zero-shot FacTax methods could be seamlessly applied with stronger LLMs to harness their ongoing development; Section 4.4 demonstrates that switching to the larger GPT-4o or Llama3-70B can unsurprisingly boost the improvement “for free”, significantly surpassing previous trained models specialized for this task.
@@
We then further seek to distill a model that fuses the task taxonomy into model parameters through supervised training. By unifying the error types of previous independently introduced datasets, we regard them jointly as training resources. Llama3-8B models are trained to learn binary decisions as well as to recognize specific error types on summaries, through our designed completions and training strategies. The resulting model outperforms previous supervised baselines, and is able to match the zero-shot FacTax performance with ChatGPT, effectively acting as an efficient alternative to zero-shot reasoning with much larger LLMs.
@@
  • We pinpoint key error types and incorporate them into our designed zero-shot prompting schemes, anchoring LLM reasoning within an explicit solution space. Experiments on diverse datasets with multiple LLMs demonstrate its efficacy.
  • We further distill a model that rationalizes the task taxonomy into parameters through our supervised training strategies, offering SOTA performance with a smaller parameter size.
@@ 2 Related Work
Factual Inconsistency Evaluation Datasets
Non-LLM Approaches
State-of-the-art models prior to LLMs mainly focus around two directions. The first is to effectively leverage NLI models to assess the entailment between the document-summary pair, such as Falke et al. (2019) and SummaC Laban et al. (2022). Within this direction, several works focus on improving the NLI model itself through methods such as synthetic data construction Kryscinski et al. (2020); Yin et al. (2021); Utama et al. (2022); Qiu et al. (2024) or multitask learning Zha et al. (2023a, b). The second direction employs QA-based models, such as QuestEval Scialom et al. (2021) and QAFactEval Fabbri et al. (2022), where they generate questions regarding explicit entities in the summary, then verify upon the source document. Apart from the two main directions, other works have also explored methods such as syntactic dependencies Goyal and Durrett (2020) or information extraction Nan et al. (2021). Particularly, Chan et al. (2023) also emphasizes interpretable factual errors by extracting semantic frames.
LLM Approaches
The capability of LLMs on detecting inconsistencies have been studied by several recent works. Most of them resolve this task in the zero-shot or few-shot prompting manner Shen et al. (2023); Luo et al. (2023); Wang et al. (2023); Liu et al. (2023a). Besides, FActScore firstly recognizes atomic claims then performs LLM inference Min et al. (2023). Other utilization of LLMs have also been proposed, such as synthetic data generation with LLMs Gekhman et al. (2023) or retrieval-augmented factuality detection Chern et al. (2023).
3 Task Taxonomy
@@
Table 1 shows the resulting conversion as well as more statistical details. We do notice that neither our adopted taxonomy nor the fine-grained one in AggreFact is completely free from noises, due to different annotation standards across datasets. Though, the current conversion is enough to bring positive gain as shown in Section 4.
@@ 4 Approach: Zero-Shot Paradigm
To incorporate the error type taxonomy, we first propose zero-shot prompting methods that leverage the general comprehension capability of LLMs, aiming to depict the explicit solution space to facilitate the zero-shot inference. As such, the proposed methods could harness the promising development of LLMs for this task, tying to greater potentials when switching to more capable LLMs.
@@ 4.1 FacTax
Our first designed prompting scheme, dubbed Factuality with Taxonomy (FacTax), follows the standard zero-shot procedure: for a document-summary pair, we instruct a LLM to determine whether the summary is factually correct, as in previous works utilizing LLMs Luo et al. (2023); Wang et al. (2023); Liu et al. (2023a). For each error type, we handcraft its explanation along with an optional example, and we ask the LLM to reason in a Chain-of-Thought (CoT) style Wei et al. (2022): whether there are any specific error types present in the summary, instead of generating a binary decision directly. A summary is thus recognized as factually correct through a rationalization stage, whenever no specified error types are present.
The resulting zero-shot inference, to this end, is regularized by the underlying task taxonomy, so to achieve a comprehensive task reasoning. We provide our full prompt in Appx. B.
@@ 4.2 FacTax-WD
After our preliminary experiments, we manually set the window size as roughly 30 words to balance between the performance and efficiency. It is worth noting that smaller window size (i.e. one sentence per window) does not necessarily lead to higher performance, as we observe that pronouns in each short window requires coreference resolution Xu and Choi (2020, 2021) that could often introduce false negatives when they are inferred without its broader context.
@@
FacTax (4o) 94.12 79.53 61.67 79.65 69.60 74.07 70.21 73.89 70.05 71.05 74.63 47.42 46.57 61.85
Table 2: Evaluation results for the zero-shot paradigm (Section 4.3). Five domains (10 datasets in total) are evaluated, where the setting for CNN/DM and XSum is kept consistent and comparable with AggreFact-FtSota Tang et al. (2023). MACRO is the final evaluation metric that computes the macro-average score across each domain. FacTax methods are our proposed approaches that ground the zero-shot inference by the task taxonomy. All LLM-based methods are shown the averaged scores of three repeated runs for robust evaluation. The same underlying LLM (gpt-3.5-turbo-0125) is adopted for direct comparison among LLM methods, except for FacTax (4o) that employs GPT-4o for demonstrating the improved performance by simply switching to stronger LLMs.
@@ 4.3 Zero-Shot Experiments
For GovReport and SQuALITY, documents are long articles that can exceed the model’s length limit. We follow Wu et al. (2023) that for each document, top sentences that maximize ROUGE scores towards the summary are retrieved as a condensed context (details in Appx. C).
@@
We adopt strong non-LLM models as baselines, including QuestEval Scialom et al. (2021), QAFactEval Fabbri et al. (2022), SummaC Laban et al. (2022), AlignScore Zha et al. (2023a) and Align Zha et al. (2023b). For each, we either take the evaluation scores from prior works, or run the code released by the authors on datasets not evaluated previously. Scores of Falsesum and AMRFact are from their original papers. For CNN/DM and XSum, we adopt thresholds per dataset to be consistent with previously reported numbers. More details on non-LLM baselines are provided in Appx. C.
@@
We use ChatGPT-ZS, ChatGPT-CoT Luo et al. (2023), ChatGPT-Star Wang et al. (2023), and G-Eval Liu et al. (2023a) as the LLM baselines, which have achieved strong performance in AggreFact. For direct comparison, we use the same ChatGPT (gpt-3.5-turbo-0125) for all LLM methods. We also vary LLMs for more insights on model comparison in Sec. 4.4.
For FacTax, we adapt our methods to additionally yield a score on GovReport and SQuALITY summaries to enable evaluation with their score-based labels.
@@ Results
Due to the variation of LLM generation, we run all LLM-based methods three times for robust conclusions, and show the averaged scores of each dataset in Table 2, along with evaluation results of non-LLM baselines. Several observations from Table 2 can be made as follows.
@@
• By switching to the stronger GPT-4o, FacTax receives significant performance boost on almost all domains (3.8% average improvement), highlighting the flexibility, simplicity and future potentials of our proposed zero-shot methods.
@@ 4.4 Zero-Shot Analysis
We focus on the three common domains: CNN/DM, XSum, Dialogues, and perform further analysis for more regarding insights.
@@
Table 3: Evaluation results using FacTax on three domains by varying LLMs. For the AggreFact-FtSota benchmark, FacTax with ChatGPT outperforms previous zero-shot G-Eval with GPT-4 (numbers reported by Qiu et al. (2024)).
@@
Apart from ChatGPT, we employ GPT-4o from OpenAI, as well as strong open-source LLMs including Llama3-8B/70B and Mistral-7B. Table 3 provides the model comparison by adopting the same FacTax prompts. Notably, FacTax with ChatGPT achieves SOTA performance on AggreFact-FtSota benchmark Tang et al. (2023), outperforming G-Eval with GPT-4 Qiu et al. (2024).
Just by switching to GPT-4o, there comes a direct boost upon ChatGPT by 3+ MACRO score overall. There is still quite a gap of almost 10% between the smaller 7B/8B models and the larger OpenAI models. By Table 3, it is evident that increasing the model size significantly improves the task reasoning, as switching Llama3-8B to 70B obtains performance gain by an impressive 12.8%, making the zero-shot method future-proof due to the rapid LLM development.
@@
Figure 2 plots the performance curve on different lengths of summaries using ChatGPT. FacTax-WD is shown more robust against long summaries, due to its length-agnostic scoring mechanism. For long summaries, false positives become more often for FacTax, which is alleviated by the window-based inference of FacTax-WD. Thus, we suggest window-based inference for long summaries (e.g. > 100 tokens), while the vanilla inference should suffice for short summaries.
@@ 5 Approach: Supervised Paradigm
As Section 4 has demonstrated the strengths of grounding zero-shot LLM inference by its task taxonomy, we further seek to distill a model that absorbs the taxonomy into LLM parameters through supervised training. Two advantages could come with such distillation. First, by learning the taxonomy from examples, the model gains real-world distribution of each error type, rather than relying on shallow comprehension through prompt instructions. Second, it is more efficient and practical compared to zero-shot methods, avoiding the need for large model sizes and lengthy generation due to the CoT reasoning.
@@ Results
Table 4 shows the evaluation results of our supervised paradigm, along with comparison by various zero-shot results. Unsurprisingly, trained Llama3-8B models of any settings outperform its zero-shot inference by large margins, up to 13 MACRO score. More importantly, I&II-Taxonomy + INF-Taxonomy achieves the best performance, matching the best existing approaches by using FacTax with GPT-4o and Llama3-70B. Our trained model can effectively serve as an efficient alternative to zero-shot inference by much larger LLMs.